# Inductive van der Waals Force between Two Quantum Loops

### v1

@@ Abstract
We study the van der Waals-London force, which is typically associated with fluctuating dipoles in atoms, in a mesoscopic circuit consisting of two inductively coupled superconducting loops. We investigate the inductive van der Waals-London interaction using both semiclassical and quantum electrodynamic (QED) approaches. The semiclassical model predicts a repulsive interaction due to anticorrelated current fluctuations. In contrast, the QED framework, which incorporates virtual photon exchange, reveals a predominantly attractive force. A key contribution comes from a state-independent two-photon exchange, which is absent in the semiclassical description and undetectable by spectroscopy. Our study introduces a new experimental platform for measuring the van der Waals force between individual artificial atoms via controlled mesoscopic circuits.
@@
where $g(R)=[\ln{(2a/R)}+\ln{4}-1]/[\ln{(2a/b)}+\ln{4}-2]^{2}$ is a geometry-dependent factor of ${\cal O}(1)$. Here, $b$ denotes the the cross-sectional radius of the loop wire. For a typical mesoscopic scale of about $R=0.1\mathrm{\mu m}$, the magnitude of the force ($\sim\hbar c/R^{2}$) is approximately $3\times 10^{-12}\mathrm{N}$. Current technology can detect this minute force (see, for example, Ref. Moser et al., 2013).
@@
Conclusion-. In summary, we have investigated the van der Waals-London interaction through inductive coupling between two superconducting quantum loops. Our study reveals a fundamental distinction between the semiclassical and quantum electrodynamic treatments of this mesoscopic system. The semiclassical approach, which is based on the shift in the zero-point energy coupled LC oscillators, predicts a repulsive interaction due to anticorrelated currents. In contrast, the QED treatment shows that the dominant contribution to the mutual force comes from the state-independent two-photon exchange term, $\Delta E_{iii}$, resulting in a predominantly attractive force with a short-distance dependence $F_{3}\sim-\hbar c/R^{2}g(R)$, reminiscent of the Casimir effect. Crucially, the key QED term is invisible in spectroscopy yet governs the mutual force ($|F_{3}/F_{12}|\gg 1$). Our findings reveal the role of quantum fluctuations in the inductive van der Waals force, specifically the $\mathbf{A}^{2}$ term in the Hamiltonian. Our work suggests new experimental avenues for observing the van der Waals force in artificial atoms via controlled mesoscopic circuits.

### v3

@@ Abstract
We study the van der Waals-London force, which is typically associated with fluctuating electric dipoles in atoms, in a mesoscopic circuit consisting of two inductively coupled superconducting loops. We investigate the inductive van der Waals-London interaction using both semiclassical and quantum electrodynamic (QED) approaches. The semiclassical model predicts a repulsive interaction due to anticorrelated current fluctuations. In contrast, the QED framework, which incorporates virtual photon exchange, reveals a predominantly attractive force. A key contribution comes from a state-independent two-photon exchange, which is absent in the semiclassical description and undetectable by spectroscopy. Our study introduces a new experimental platform for measuring the van der Waals force between individual artificial atoms via controlled mesoscopic circuits.
@@
where $g(R)=[\ln{(2a/R)}+\ln{4}-1]/[\ln{(2a/b)}+\ln{4}-2]^{2}$ is a geometry-dependent factor of ${\cal O}(1)$. Here, $b$ denotes the cross-sectional radius of the loop wire. For a typical mesoscopic scale of about $R=0.1\mathrm{\mu m}$, the magnitude of the force ($\sim\hbar c/R^{2}$) is approximately $3\times 10^{-12}\mathrm{N}$. Current technology can detect this minute force (see, for example, Ref. Moser et al., 2013).
@@
Conclusion-. In summary, we have investigated the van der Waals-London interaction through inductive coupling between two superconducting quantum loops. Our study reveals a fundamental distinction between the semiclassical and quantum electrodynamic treatments of this mesoscopic system. The semiclassical approach, based on the shift in the zero-point energy in coupled LC oscillators, predicts a repulsive interaction due to anticorrelated currents. In contrast, the QED treatment shows the opposite result. The dominant contribution to the mutual force comes from the state-independent two-photon exchange term, $\Delta E_{iii}$, resulting in a predominantly attractive force with a short-distance dependence $F_{3}\sim-\hbar c/R^{2}g(R)$, reminiscent of the Casimir effect. The key QED term is crucial yet invisible in spectroscopy, governing the mutual force ($|F_{3}/F_{12}|\gg 1$). The role of vacuum fluctuations in QED has been highlighted through various interesting phenomena such as the Lamb shift, spontaneous emission, and others. Our finding is another interesting manifestation of this and reveals the role of quantum fluctuations in the inductive van der Waals force, specifically the $\mathbf{A}^{2}$ term in the Hamiltonian. The QED-specific term has a striking effect on the interaction, turning the repulsive force into an attractive one. Our work suggests new experimental avenues for observing the van der Waals force in artificial atoms via controlled mesoscopic circuits. Furthermore, this could be a new kind of experimental test for the validity of QED.